In 2018, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2020. Information related to the contract is as follows:

                                     2018                         2019                            2020

Cost incurred during the year    $ 2,291,000 $ 3,555,000 $ 2,259,400

Estimated costs to complete as of year-end         5,609,000 2,054,000 0

Billings during the year    2,110,000               3,736,000                   4,154,000

Cash collections during the year                            1,855,000    3,500,000 4,645,000

Westgate recognizes revenue over time according to percentage of completion. rev: 09_15_2017_QC_CS-99734 Required:

1. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years

                                          2018   2019     2020

Revenue:

Gross profit:
Answer #1

2018

2019

2020

Contract price

10,000,000

10,000,000

10,000,000

Actual costs to date

2,291,000

5,846,000

8,105,400

Estimated costs to complete

5,609,000

2,054,000

0

Total estimated costs

7,900,000

7,900,000

8,105,400

Estimated gross profit (loss) {actual in 2020}

2,100,000

2,100,000

1,894,600

Revenue recognition using

= {actual costs to date/total estimated costs * contract price} – revenue recognized till            date

2018 = 2291000/7900000 * 10,000,000 = $2,900,000

2019 = {5846000/7900000 * 10,000,000} – 2900,000 = $7,400,000 -2,900,000 = $4,500,000

2020 = 10,000,000 – (2900,000 + 4500,000) = $2,600,000

Similarly, Gross profit (loss) recognition using percentage of completion method as

2018 = 2291000/7900000 * 2,100,000 = $609,000

2019 = {5846000/7900000 * 2100,000} – 609000 = 1554000 – 609000 = $945000

2020 = 1894600 – 1554000 = $340,600